OpenStudy (anonymous):

The square of x is equal to 4 times the square of y. If x is 1 more than twice y, what is the value of x? i) -4 ii) - 1/2 iii) - 1/4 iv) 1/4 v) 1/2

OpenStudy (anonymous):

\[x^2=4y^2\] \[x=2y+1\] Use substitution.

OpenStudy (aravindg):

@amishra try the method suggested above^

OpenStudy (anonymous):

I tried, but I am not getting the correct answer :|

OpenStudy (anonymous):

Okay...Everytime you see the letter x, substitute this: (2y+1)

OpenStudy (anonymous):

\[x^2=4y^2\] \[(2y+1)^2=4y^2\]

OpenStudy (anonymous):

Now you can do it using your knowledge of expanding.

OpenStudy (anonymous):

from the eq. x=2y+1 we get, y=(x-1)/2 now x^2=4y^2 =>x^2=4[(x-1)/2]^2=4[(x-1)^2]/4=(x-1)^2=x^2-2x+1 =>2x-1=0 =>x=1/2

OpenStudy (anonymous):

@abhrag321 Good job going against the Code of Contact of sharing answers and not actually helping them understand the question.

OpenStudy (anonymous):

Code of Conduct*

OpenStudy (anonymous):

Its okay. I finally understood my mistake. I was searching for y instead of x :P THANK YOU SO MUCH FOR YOUR HELP!!!!!!!!!!

OpenStudy (anonymous):

I have read the Code of Conduct. There is the line, "Don't provide someone with just the answer - explain the process, and help guide them through understanding the problem." So please explain the violation process.